# Flow Matching and Diffusion Models via PointNet for Generating Fluid Fields on Irregular Geometries

### v1

@@ 4.6 Robustness
The objective of this subsection is to evaluate how robust the three deep learning frameworks, Flow Matching PointNet, Diffusion PointNet, and baseline PointNet, are in handling such incomplete data. Table 4, together with Figures 7 and 8, illustrate the results of this analysis. When 5%, 10%, and 15% of the points in each point cloud (e.g., 15% of 1024 points) are randomly removed, the corresponding errors in the $x$- and $y$-components of the velocity vector (i.e., $u$ and $v$) and in the pressure variable (i.e., $p$) are listed in Table 4. As expected, increasing the percentage of missing points leads to higher prediction errors. Among the three models considered, the Flow Matching PointNet demonstrates the highest robustness, followed by the Diffusion PointNet, while the baseline PointNet exhibits the lowest robustness. When 10% of the points are removed, the average relative pointwise error in the $L^{2}$ norm for $v$ is approximately 6.9% for Flow Matching PointNet, 8.6% for Diffusion PointNet, and 14.9% for baseline PointNet. Notably, the error for baseline PointNet exceeds 10%, which is generally considered unacceptable in engineering design contexts. When 15% of the points are removed, the prediction error for the $v$ increases to about 8.3% for Flow Matching PointNet, nearly 11% for Diffusion PointNet, and over 17% for baseline PointNet. Under these conditions, Flow Matching PointNet is the only model maintaining an error below 10%.
Figures 7 and 8 show the predictions obtained by the Flow Matching PointNet and Diffusion PointNet, respectively, for the $u$, $v$, and $p$ fields when 15% of the points are removed. It is important to note that some of the missing points are located on the surface of the cylinder. Mathematically, the velocity and pressure fields are functions of the cylinder’s cross-sectional geometry. Consequently, the learning process of the deep network inherently depends on this geometry. Kashefi et al. [29] demonstrated that such boundary points are, in fact, critical within the PointNet framework and play an essential role in geometric deep learning. This finding further highlights the importance of robustness and underscores the superior performance of the Flow Matching PointNet in handling geometries with missing boundary points.
@@ 5 Summary and outlook
In this study, we introduced, for the first time, the Flow Matching PointNet and Diffusion PointNet, which integrate the formulations of flow matching [1] and diffusion models [2] into PointNet [34, 29, 30] to generate velocity and pressure fields of fluid flows conditioned on the geometry of the computational domain. In these generative frameworks, irregular geometries were represented as point clouds, where each point corresponded to a spatial coordinate. Three noise vectors, corresponding to the $x$-velocity component, the $y$-velocity component, and the pressure, were concatenated with the spatial position vectors $x$ and $y$. In addition, sinusoidal positional encoding vectors were concatenated to these to form the input features. A forward processing phase was then initiated, during which the Diffusion PointNet was trained to predict the noise components added to the velocity and pressure fields, while the Flow Matching PointNet, in a similar manner, predicted the derivatives of the noise with respect to the temporal variable. After training, during inference, a reverse processing phase was conducted, where noise conditioned on an unseen geometry was transformed to generate the velocity and pressure fields corresponding to that geometry. In the Flow Matching PointNet, we used the explicit Euler numerical scheme during the reverse process.
@@
To demonstrate the effectiveness of the proposed frameworks, the classical problem of fluid flow around a cylinder was employed as a benchmark. In this setup, the cross-section of the cylinder varied across different geometries, resulting in corresponding variations in the pressure and velocity distributions. These variations constituted the dataset that was used for training. The analysis of the predictions from the Flow Matching PointNet and Diffusion PointNet models revealed that their performance was significantly superior to that of the baseline PointNet framework. In particular, employing the Flow Matching PointNet and Diffusion PointNet models reduced the average relative pointwise error in the $L^{2}$ norm to below 10%, whereas the error for the baseline PointNet model exceeded 10%. Flow Matching PointNet and Diffusion PointNet demonstrated improved accuracy in predicting pressure-induced lift and drag forces. Furthermore, since the proposed frameworks generate predictions based on stochastic noise inputs, multiple realizations can be produced for any given geometry, enabling estimation of the uncertainty in model predictions. This capability is especially valuable for engineering design and safety factor considerations. Specifically, we demonstrated the uncertainty range associated with the maximum possible errors in pressure-induced drag and lift predictions. Additionally, we investigated the robustness of the Flow Matching PointNet and Diffusion PointNet models in cases where the geometry was incomplete (i.e., when a percentage of points in a point cloud was missing). The results showed that both models exhibited greater robustness compared to the baseline PointNet, with the Diffusion PointNet capable of maintaining the average relative pointwise error in the $L^{2}$ norm of the velocity and pressure fields below 10% even when up to 15% of the point cloud data were missing. Finally, although both the Flow Matching PointNet and Diffusion PointNet significantly outperformed the baseline PointNet, the overall difference between the two was not substantial. The Diffusion PointNet performed slightly better in predicting pressure, leading to more accurate pressure-induced drag and lift computations, whereas the Flow Matching PointNet demonstrated greater robustness when the geometry was partially missing.

### v2

@@ 4.6 Robustness
The objective of this subsection is to evaluate how robust the three deep learning frameworks, Flow Matching PointNet, Diffusion PointNet, and baseline PointNet, are in handling such incomplete data. Table 4, together with Figures 18 and 19, illustrate the results of this analysis. When 5%, 10%, and 15% of the points in each point cloud (e.g., 15% of 1024 points) are randomly removed, the corresponding errors in the $x$- and $y$-components of the velocity vector (i.e., $u$ and $v$) and in the pressure variable (i.e., $p$) are listed in Table 4. As expected, increasing the percentage of missing points leads to higher prediction errors. Among the three models considered, the Flow Matching PointNet demonstrates the highest robustness, followed by the Diffusion PointNet, while the baseline PointNet exhibits the lowest robustness. When 10% of the points are removed, the average relative pointwise error ($L^{2}$ norm) for predicting the $v$ variable is approximately 6.9% for Flow Matching PointNet, 8.6% for Diffusion PointNet, and 14.9% for baseline PointNet. Notably, the error for baseline PointNet exceeds 10%, which is generally considered unacceptable in engineering design contexts. When 15% of the points are removed, the prediction error for the $v$ increases to about 8.3% for Flow Matching PointNet, nearly 11% for Diffusion PointNet, and over 17% for baseline PointNet. Under these conditions, Flow Matching PointNet is the only model maintaining an error below 10%.
Figures 18 and 19 show, respectively, the predictions obtained by the Flow Matching PointNet and Diffusion PointNet for the $u$, $v$, and $p$ variables when 15% of the points are removed. It is important to note that some of the missing points are located on the surface of the cylinder. Mathematically, the velocity and pressure fields are functions of the cylinder’s cross-sectional geometry. Consequently, the learning process of the deep network inherently depends on this geometry. Kashefi et al. [29] demonstrated that such boundary points are, in fact, critical within the PointNet framework and play an essential role in geometric deep learning. This finding further highlights the importance of robustness and underscores the superior performance of the Flow Matching PointNet in handling geometries with missing boundary points.
@@ 5 Summary and outlook
In this study, we introduced, for the first time, the Flow Matching PointNet and Diffusion PointNet, which integrated the formulations of flow matching [1] and diffusion models [2] into PointNet [34, 29, 30] to generate velocity and pressure fields of fluid flows conditioned on the geometry of the computational domain. In these generative frameworks, irregular geometries were represented as point clouds. Independent standard Gaussian noise fields, corresponding to the velocity components in the $x$- and $y$-directions as well as the pressure field, were concatenated with the spatial coordinates. In addition, sinusoidal positional encodings were concatenated with the spatial coordinates and noise fields to form the input features. A forward processing phase was then initiated, during which the Diffusion PointNet was trained to predict the noise components added to the velocity and pressure fields, while the Flow Matching PointNet, in a similar manner, predicted the derivatives of the noise with respect to a temporal variable. After training, inference was performed via a reverse processing phase, in which independent standard Gaussian noise fields associated with the velocity components and pressure were conditioned on an unseen geometry to generate the corresponding velocity and pressure fields. In Flow Matching PointNet, we used the explicit Euler numerical scheme during the reverse process.
@@
To demonstrate the effectiveness of the proposed frameworks, the classical problem of fluid flow around a cylinder was employed as a benchmark. In this setup, the cross-section of the cylinder varied across different geometries, resulting in corresponding variations in the velocity and pressure distributions. These variations constituted the dataset that was used for training. The analysis of the predictions from the Flow Matching PointNet and Diffusion PointNet models revealed that their performance was significantly superior to that of the baseline PointNet framework. In particular, employing the Flow Matching PointNet and Diffusion PointNet models reduced the average relative pointwise error in the $L^{2}$ norm to below 10%, whereas the error for the baseline PointNet model exceeded 10%. Flow Matching PointNet and Diffusion PointNet demonstrated improved accuracy in predicting pressure-induced lift and drag forces. Furthermore, since the proposed frameworks generated predictions based on standard Gaussian noise inputs, multiple realizations were produced for any given geometry, enabling estimation of the uncertainty in model predictions. This capability was especially valuable for engineering design and safety factor considerations. Specifically, we demonstrated the uncertainty range associated with the maximum possible errors in pressure-induced drag and lift predictions. Additionally, we investigated the robustness of the Flow Matching PointNet and Diffusion PointNet models in cases where the geometry was incomplete (i.e., when a percentage of points in a point cloud was missing). The results showed that both models exhibited greater robustness compared to the baseline PointNet, with the Diffusion PointNet capable of maintaining the average relative pointwise error in the $L^{2}$ norm of the velocity and pressure fields below 10% even when up to 15% of the point cloud data were missing. Finally, although both the Flow Matching PointNet and Diffusion PointNet significantly outperformed the baseline PointNet, the overall difference between the two was not substantial. The Diffusion PointNet performed slightly better in predicting pressure, leading to more accurate pressure-induced drag and lift computations, whereas the Flow Matching PointNet demonstrated greater robustness when the geometry was partially missing.